Members of the Iroquois nation played a game in which there were 6 flat disks which had different colours on each side (say red and blue). People would take turns throwing the disks in the air and noting how they landed. If they were all the same colour, the person scored 3 points. If all but one were the same colour the person would score 1 point. For any other outcome, no points were assigned. The winner was the first to 20 points (or more).
A) calculate the probability of all the possible outcomes grouped together (for example, one grouping would be all outcomes in which there are 4 blue and 2 red).
I've tried: P(E)= 1/6P2 and other variations like that, I just keep getting smaller numbers and don't think I'm doing it correctly. Please help.

Answer by ewatrrr(24785)

p(blue) = 1/2, p(red) = 1/2, n = 6
People would take turns throwing the disks
P(all same color) = (1/2)^6
P(5blue 0r 5red) = 6C5(1/2)^5(1/2)^1 + 6C5(1/2)^5(1/2)^1
etc...